The plasma density in the lower ionosphere has been measured during satellite re-entry to be about 10^18 m^-3 at 50km altitude,10^17at 70 km,and 10^14 at 8 km.what are the plasma frequencies there?...... please solve it completely
1
Expert's answer
2020-06-17T09:29:23-0400

The plasma frequency, depending on the number density n, charge of electron e, and mass m is


"f=\frac{1}{2\pi}\sqrt{\frac{ne^2}{m\epsilon_0}}."


At 8 km:


"f=224\cdot10^{15}\text{ Hz},"

at 50 km:


"f=22.4\cdot10^{18}\text{ Hz},"

at 70 km:


"f=7.09\cdot10^{15}\text{ Hz}."